The width of a rectangle is 1 inch more then twice it's length, and the perimeter is is 104 inches. Find the length and width of the rectangle.
Answer by waynest(281)

width = 2L + 1
length = L
p = 104
2L + 2(2L + 1) = 104
2L + 4L + 2 = 104
6L + 2 - 2 = 104 - 2
6L = 102
6L/6 = 102/6
L = 17
check:
2(17) + 2(2(17) + 1) = 104
34 + 2(34 + 1) = 104
34 + 68 + 2 = 104
104 = 104